OpenStudy (anonymous):

Two planes which are 1820 miles apart, fly toward each other. Their speeds differ by 60 mph. If they pass each other in 2 hours, what is the speed of each?

OpenStudy (anonymous):

d= rt 2r + 2(r+60) = 1820

OpenStudy (anonymous):

4r + 120 = 1820 4r = 1820 -120

OpenStudy (anonymous):

r = 425

OpenStudy (anonymous):

and second plane has 485

OpenStudy (anonymous):

Alright, lets get some equations down using that data: The distance each plane is going to fly is going to be its velocity times 2 hours (they will both be flying for 2 hours. So we have: \[d_1=v_1(2), d_2=v_2(2)\] Let v1 be the velocity of the slower plane. Then v2 is: \[v_2 = v_1+60\] also, their distances must add up to 1820, so now we have: \[1820 = v_1(2)+(v_1+60)(2) \Rightarrow v_1 = \frac{1820-120}{4} = 425mph\] That is the slower plane, add 60 to get the faster plane going at 485 mph.